1. A company developed a promising product and wants to make a decision. Based on the best available information, a decision tree was drawn.

Payoff $ millions) (.5) Large Market Produce & Market (.5) Low Mkt. 2.0 (.5) Large Market Develop Sell Idea 3.0 Product (.5) Low Mkt. 1.0 Lease for Royalty (.5) Large Market 2.2 Sell Idea CompanyA (.5) Low Mkt. 1.2 (.5) Large Market Company EB 1.9 (.5) Low Mkt. 1.3 1.0

a) Calculate expected value at every node.

b) If they “develop the product” & “produce & market”, how much can it expect?

c) What is the optimum decision?   
Answer #1

a) In order to calculate the expected value we need to multiply the payoff against the probability. The expected values are calculated only for the chance node (round nodes).

Develop product

Produce and market

0.5*5 + 0.5*-2 = 1.5

Sell Idea

0.5*3 + 0.5*1 = 2

Lease for royalty

Company A

0.5*2.2 + 0.5*1.2 = 1.7

Company B

0.5*1.9 + 0.5*1.3 = 1.6

b) If they develop product and produce and market, they can expect (the first node)

0.5*5 + 0.5*-2 = 1.5 million

c) The best outcome of develop product is 2 million by selling the idea.

The best outcome of lease for royalty is 1.7 million by company A

Selling the idea provides 1 million directly

The optimum decision is to develop the product and then sell the idea. This will give a expected payoff of 2 million.